# Comparison of lubrication theory and Stokes flow in corner geometries with flow separation

### v3

@@ 2 Lubrication Theory
(1) $\frac{\partial p}{\partial x}$ $=\eta\Big(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\Big)-\rho\Big(u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}\Big),$
(2) $\frac{\partial p}{\partial y}$ $=\eta\Big(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\Big)-\rho\Big(u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}\Big),$
@@
We introduce the characteristic length scales $L_{x}=x_{f}-x_{0}$ and $L_{y}=\max h(x)>0$, and the length scale ratio $\varepsilon=L_{y}/L_{x}$. For a prescribed constant flux $\mathcal{Q}\neq 0$, define the characteristic velocities $U_{*}=\mathcal{Q}/L_{y}$ and $V_{*}=\mathcal{Q}/L_{x}$. Or, in the case of $\mathcal{Q}=0$ and $\mathcal{U}\neq 0$, let $U_{*}=\mathcal{U}$ and $V_{*}=U_{*}L_{y}/L_{x}$. The Reynolds number is given by $\text{Re}=\rho U_{*}{\color[rgb]{0,0,0}L_{x}}/\eta$, and the characteristic pressure is $P_{*}=\eta U_{*}L_{x}/L_{y}^{2}$. Upon nondimensionalization and applying the lubrication assumptions $\varepsilon^{2}\ll 1$ and $\varepsilon^{2}\text{Re}\ll 1$, the incompressible Navier-Stokes equations Eqs. 1 and 2 are reduced to the governing equations of lubrication theory:
@@ 3 The Stokes equation
The Stokes equation is obtained from the incompressible Navier-Stokes equations in the limit of zero Reynolds number [leal_advanced_2007]. With the standard dimensionless variables, $X=x/L_{x}$, $Y=y/L_{x}$, $U=u/U_{*}$, $V=v/U_{*}$, and $P=p/P_{*}$, the Navier-Stokes equations Eqs. 1 and 2 are expressed in dimensionless terms as,
(15) $\frac{\partial P}{\partial X}$ $=\Big(\frac{\partial^{2}U}{\partial X^{2}}+\frac{\partial^{2}U}{\partial Y^{2}}\Big)-\text{Re}\Big(U\frac{\partial U}{\partial X}+V\frac{\partial U}{\partial Y}\Big),$
(16) $\frac{\partial P}{\partial Y}$ $=\Big(\frac{\partial^{2}V}{\partial X^{2}}+\frac{\partial^{2}V}{\partial Y^{2}}\Big)-\text{Re}\Big(U\frac{\partial V}{\partial X}+V\frac{\partial V}{\partial Y}\Big).$
Compared with lubrication theory, this nondimensionalization is performed using a single length scale $L_{x}$ and velocity scale $U_{*}$. Furthermore, the characteristic pressure is now $P_{*}=\eta U_{*}/L_{x}$. Consequently, Eqs. 15 and 16 are not sensitive to the length scale ratio $\varepsilon$ that appears in lubrication theory.
@@
the dimensionless Navier-Stokes equations Eqs. 15 and 16 may be expressed in stream-velocity form,
@@ Appendix B Convergence
In the absence of an exact solution to the Stokes equation, we measure the grid convergence of the numerical solution by comparing to the solution on the finest grid resolution available. We denote the grid scale $\mathcal{N}=\frac{1}{\Delta x}=\frac{1}{\Delta y}$. Convergence in the stream function $\psi$ for the BFS with $\mathcal{H}=2$ and for the triangular cavity with $H=4$ are shown in Fig. 21. For the BFS, the convergence is order $\mathcal{O}(\Delta x^{2})$. For the triangular cavity, the convergence in $\psi$ is between order $\mathcal{O}(\Delta x^{2})$ and $\mathcal{O}(\Delta x)$. We can expect a reduced order of convergence in the triangular cavity at $\mathcal{H}=4$ due to the boundary smoothing in the acute corners of the triangle.

### v4

@@ 2 Lubrication Theory
(1) $\frac{\partial p}{\partial x}$ $=\eta\Big(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\Big)-\rho\Big(\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}\Big),$
(2) $\frac{\partial p}{\partial y}$ $=\eta\Big(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\Big)-\rho\Big(\frac{\partial v}{\partial t}+u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}\Big),$
@@
We introduce the characteristic length scales $L_{x}=x_{f}-x_{0}$ and $L_{y}=\max h(x)>0$, and the length scale ratio $\varepsilon=L_{y}/L_{x}$. For a prescribed constant flux $\mathcal{Q}\neq 0$, define the characteristic velocities $U_{*}=\mathcal{Q}/L_{y}$ and $V_{*}=\mathcal{Q}/L_{x}$. Or, in the case of $\mathcal{Q}=0$ and $\mathcal{U}\neq 0$, let $U_{*}=\mathcal{U}$ and $V_{*}=U_{*}L_{y}/L_{x}$. The Reynolds number is given by $\text{Re}=\rho U_{*}{\color[rgb]{0,0,0}L_{x}}/\eta$, the characteristic time is $T_{*}=L_{x}/U_{*}$, and the characteristic pressure is $P_{*}=\eta U_{*}L_{x}/L_{y}^{2}$. Upon nondimensionalization and applying the lubrication assumptions $\varepsilon^{2}\ll 1$ and $\varepsilon^{2}\text{Re}\ll 1$, the incompressible Navier-Stokes equations Eqs. 1 and 2 are reduced to the governing equations of lubrication theory:
@@ 3 The Stokes equation
The Stokes equation is obtained from the incompressible Navier-Stokes equations in the limit of zero Reynolds number [leal_advanced_2007]. With the standard dimensionless variables, $X=x/L_{x}$, $Y=y/L_{x}$, $U=u/U_{*}$, $V=v/U_{*}$, $T=t/T_{*}$ and $P=p/P_{*}$, the Navier-Stokes equations Eqs. 1 and 2 are expressed in dimensionless terms as,
(15) $\frac{\partial P}{\partial X}$ $=\Big(\frac{\partial^{2}U}{\partial X^{2}}+\frac{\partial^{2}U}{\partial Y^{2}}\Big)-\text{Re}\Big(\frac{\partial U}{\partial T}+U\frac{\partial U}{\partial X}+V\frac{\partial U}{\partial Y}\Big),$
(16) $\frac{\partial P}{\partial Y}$ $=\Big(\frac{\partial^{2}V}{\partial X^{2}}+\frac{\partial^{2}V}{\partial Y^{2}}\Big)-\text{Re}\Big(\frac{\partial V}{\partial T}+U\frac{\partial V}{\partial X}+V\frac{\partial V}{\partial Y}\Big).$
Compared with lubrication theory, this nondimensionalization is performed using a single length scale $L_{x}$ and velocity scale $U_{*}$. The characteristic pressure is now $P_{*}=\eta U_{*}/L_{x}$. Consequently, Eqs. 15 and 16 are not sensitive to the length scale ratio $\varepsilon$ that appears in lubrication theory.
@@
and assuming a steady state flow, the dimensionless Navier-Stokes equations Eqs. 15 and 16 may be expressed in stream-velocity form,
@@ Appendix B Convergence
In the absence of an exact solution to the Stokes equation, we measure the grid convergence of the numerical solution by comparing to the solution on the finest grid resolution available. We denote the grid scale $\mathcal{N}=\frac{1}{\Delta x}=\frac{1}{\Delta y}$. Convergence in the stream function $\psi$ for the BFS with $\mathcal{H}=2$ and for the triangular cavity with $H=4$ are shown in Fig. 21. For the BFS, the convergence is order $\mathcal{O}(\Delta x^{2})$. For the triangular cavity, the convergence in $\psi$ is between order $\mathcal{O}(\Delta x^{2})$ and $\mathcal{O}(\Delta x)$. We expect a reduced order of convergence in the triangular cavity at $\mathcal{H}=4$ due to the boundary smoothing in the acute corners of the triangle.